Product Cost Method of Product Costing

Voice Com, Inc., uses the product cost method of applying the cost-plus approach to product pricing. The costs of producing and selling 5,090 units of cell phones are as follows:

Variable costs: Fixed costs:
Direct materials $62 per unit Factory overhead $200,500
Direct labor 30 Selling and admin. exp. 71,200
Factory overhead 23
Selling and admin. exp. 19
Total variable cost per unit $134 per unit

Voice Com desires a profit equal to a 14% rate of return on invested assets of $601,100.

a. Determine the amount of desired profit from the production and sale of 5,090 units of cell phones.
$

b. Determine the product cost per unit for the production of 5,090 of cell phones. If required, round your answer to nearest dollar.
$ per unit

c. Determine the product cost markup percentage (rounded to two decimal places) for cell phones.
%

d. Determine the selling price of cell phones. Round to the nearest dollar.

Total Cost $per unit
Markup per unit
Selling price $per unit
Answer #1

Solution a:

Amount of desired profit = $601,100*14% = $84,154

Solution b:

Product cost per unit = Direct material + Direct labor + Variable factory overhead + Fixed Factory overhead

= $62 + 30+ 23 + ($200500/5090) = $154.3909 = $154 (Rounded)

Solution c:

Total Markup required = Desired profit + Selling and adm. expenses = $84154 + (5090*$19) + $71200 = $252,064

Total Product Cost = 5090*(62+30+23) + $200500 = $785,850

Markup percentage = $252064/ $785850 = 32.08%

Solution d:

Total cost = $154 per unit

Markup = ($252064/5090) = $50 per unit

Selling price = $154 + $50 = $204 per unit